OpenStudy (anonymous):

x + y = 6 -x + 3y = -2

OpenStudy (anonymous):

is there any 2nd euattion

OpenStudy (anonymous):

We need a second equation.

OpenStudy (anonymous):

to solve it simultaneously

OpenStudy (anonymous):

The second equation is there, @UditKulka

OpenStudy (anonymous):

adding the 2 euations we get

OpenStudy (anonymous):

y=1

OpenStudy (anonymous):

don't just say it, make her/him see how you did it too

OpenStudy (anonymous):

x=1

OpenStudy (anonymous):

k

OpenStudy (anonymous):

-8x+8x=0

OpenStudy (anonymous):

3y-2y=y

OpenStudy (anonymous):

-5+6=1

OpenStudy (anonymous):

therefore y=1

OpenStudy (anonymous):

substituting y in 8x - 2y = 6 we get

OpenStudy (anonymous):

8x - 2(1) = 6

OpenStudy (anonymous):

6/-2=8x

OpenStudy (anonymous):

-3/8=x

OpenStudy (anonymous):

i was wrong before but i got it right

OpenStudy (anonymous):

cool i have another question n thanks Uditkulka

OpenStudy (anonymous):

say them :)

OpenStudy (anonymous):

k

OpenStudy (anonymous):

x + y = 3 2x - y = 3

OpenStudy (anonymous):

Add the equations

OpenStudy (anonymous):

y-y=0

OpenStudy (anonymous):

x+2x=3x

OpenStudy (anonymous):

3x=6

OpenStudy (anonymous):

x=6/3

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

substitue x in x + y = 3

OpenStudy (anonymous):

2 + y =3 y= 3-2 y=1

OpenStudy (anonymous):

@9ust_Luv1:Got it ?

OpenStudy (anonymous):

wow you r really smart and thanks and i have it yall are a big help

OpenStudy (anonymous):

any time!

OpenStudy (anonymous):

x + y = 6 -x + 3y = -2

OpenStudy (anonymous):

x cancels out

OpenStudy (anonymous):

4y=-4

OpenStudy (anonymous):

y=-1

OpenStudy (anonymous):

x -1 =6 x=6+1 x=7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

last one

OpenStudy (anonymous):

x + y = 8 2x = y - 5

OpenStudy (anonymous):

take the y on other side

OpenStudy (anonymous):

y-y =0

OpenStudy (anonymous):

x +2x =3x

OpenStudy (anonymous):

3x =3 x=3/3 x=1

OpenStudy (anonymous):

1+y =8

OpenStudy (anonymous):

8-1

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

k

OpenStudy (anonymous):

UditKulka is correct. Using the elimination method then subbing the variable you found back into either of the equations will find the other variable. You can also check your answers by plugging n the variables you solved for into either equation and if they come out equal, then they are correct.

OpenStudy (anonymous):

oh okay thanks kelliegirl33

OpenStudy (anonymous):

no problem....I didn't really do anything...UditKulka did it all

OpenStudy (anonymous):

iik